OpenStudy (anonymous):

PLEASE HELP ME !! The 1997 a value of an object was $5000. In 2012 , it was worth $9500. The annual percent growth has been constant. What is the annual percent growth?? A) 1.37% B) 2.37% C) 3.37% D ) 4. 37%

OpenStudy (amistre64):

1997: 5000 1998: 5000(1+r) 1999: 5000(1+r)(1+r) 2000: 5000(1+r)(1+r)(1+r) 1997+n = 5000(1+r)^n so we need to determine how many years have passed then apply A = P(1+r)^n A/P = (1+r)^n nrt(A/P) = 1+r nrt(A/P) -1 = r

OpenStudy (anonymous):

Im lost ..

OpenStudy (anonymous):

I know that between 1997 - 2012 , passed 15 years , & that $4,500 increased between those years

OpenStudy (amistre64):

then we just fill in the values (9500/5000)^(1/15) -1 = r

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%289500%2F5000%29%5E%281%2F15%29+-1 multiply by 100 to get it to look like an option

OpenStudy (anonymous):

oh I got d ??